In genetics, single nucleotide polymorphisms (SNPs) are locations in the (human) genome that exhibit variation across the population. SNPs cause the differences we see in traits such as hair color. Each SNP typically has two possible alleles say A and a and each persons genotype at the SNP is either AA, Aa, or aa, where one allele comes from the persons mother and one from the father. Let X be the number of A alleles at a particular SNP, and suppose we collect a random sample of people from some population. Under some assumptions (such as random mating and no selection) we may assume that where p is called the allele frequency of allele A. What is the maximum likelihood estimator of p? What is the maximum likelihood estimate of the allele frequency of allele A if our sample consists of five people with genotypes AA, aa, Aa, aa, Aa at this particular SNP?
Answer #1

The probability mass function of Xi ~ Binom(2, p) is,

f(x_i;p) = inom{2}{x_i} p^{x_i} (1-p)^{2-x_i}

The likelihood function of X = (X1, X2, ..., Xn) is,

f(x;p) = inom{2}{x_1} p^{x_1} (1-p)^{2-x_1} imes inom{2}{x_2} p^{x_2} (1-p)^{2-x_2} .... imes inom{2}{x_n} p^{x_n} (1-p)^{2-x_n}

f(x;p) = prod_{i=1}^{n} inom{2}{x_i} p^{x_i} (1-p)^{2-x_i}

The log-likelihood function is,

inf(x: P) _ Σ In ( + Σ.rilnp + Σ(2-x.)ln(1-P) i=1 i=1

For maximum likelihood estimate,

rac{partial }{partial p}ln f(x;p) = 0

plnf(x; p) Σ.ri/p _ Σ 2-r.)/(1-p)-0

i= 1 i= 1

i= 1 i= 1

Rightarrow 1/p - 1 = 2n/sum_{i=1}^{n}x_i-1

Rightarrow 1/p = 2n/sum_{i=1}^{n}x_i

Rightarrow p = sum_{i=1}^{n}x_i /2n

Thus, the maximum likelihood estimate of the allele frequency is,

hat{p} = sum_{i=1}^{n}x_i /2n

For a given sample,

AA, aa, Aa, aa, Aa

n = 5 and the frequency of the allele in each of the five people are X = (2, 0, 1, 0, 1) . We get this by counting number of A alleles in each genotype.

sum_{i=1}^{n}x_i = 2 + 0 + 1 + 0 + 1 = 4

The maximum likelihood estimate of the allele frequency is,

hat{p} = sum_{i=1}^{n}x_i /2n = 4 / (2 * 5) = 2/5 = 0.4